# A Hessian for Gaussian Mixture Likelihoods in Nonlinear Least Squares

### v1

@@ V Proposed Approach: Hessian-Sum-Mixture
Rather, the proposed Hessian-Sum-Mixture Hessian approximation is derived by setting the second derivatives of the component errors to zero, and using the chain rule to take into account the LogSumExp nonlinearity and derive the cost Hessian.
@@ VI Nonlinear Least Squares Compatibility
and substituting into the GN Hessian approximation (6) results in exactly the Hessian-Sum-Mixture Hessian (27) as well as the loss Jacobian (26). However, the resulting cost function $\left\|\mbf{e}_{\textrm{solver},1}\right\|_{2}^{2}$ is not of the same form as (12), which can cause issues with methods that use the cost function to guide the optimization such as Levenberg-Marquardt (LM) [17]. To rectify this, an additional error term is required to make the loss into the same form as (12). The proposed error and Jacobian to be input into the solver are thus
$\mbf{e}_{\textrm{solver}}^{\mathsf{T}}$ $=\begin{bmatrix}\mbf{e}_{\textrm{solver},1}^{\mathsf{T}}&\sqrt{{\gamma_{\text{{{HSM}}}}}+\Delta J}\end{bmatrix},$ (37)
$\mbf{J}_{\textrm{solver}}^{\mathsf{T}}$ $=\begin{bmatrix}\mbf{J}_{\textrm{solver, 1}}^{\mathsf{T}}&\mbf{0}\end{bmatrix},$ (39)
where $\Delta J$ is the difference between the desired loss $J_{\textrm{GMM}}$ (12) and the norm of $\mbf{e}_{\textrm{solver},1}$,
$\Delta J=J_{\textrm{GMM}}-\left\|\mbf{e}_{\textrm{solver},1}\right\|_{2}^{2},$ (40)
and ${\gamma_{\text{{{HSM}}}}}$ is a normalization constant that ensures positivity of the square root in (37). The evaluated norm of $\mbf{e}_{\textrm{solver}}$ is thus
$\left\|\mbf{e}_{\textrm{solver}}\right\|_{2}^{2}$ $=J_{\textrm{GMM}}+{\gamma_{\text{{{HSM}}}}}.$ (41)
Finding a lower bound on $\Delta J$ that may be used as the normalization constant is difficult. The proposed solution is to set ${\gamma_{\text{{{HSM}}}}}$ to a small initial value. Each time the error is evaluated, the solver determines if ${\gamma_{\text{{{HSM}}}}}+\Delta J$ is less than zero. If yes, ${\gamma_{\text{{{HSM}}}}}$ is updated as ${\gamma_{\text{{{HSM}}}}}\leftarrow\left|\Delta J\right|+\Delta{\gamma_{\text{{{HSM}}}}}$, where $\Delta{\gamma_{\text{{{HSM}}}}}$ is a small positive value. This amounts to moving the entire cost surface upwards. While heuristic, this approach works well in practice and gives almost indistinguishable convergence from a version of the solver that uses the original loss and specifies the Hessian itself.
@@ VII Simulation and Experiments
To show the benefits of the proposed approach, the performance of the Max-Mixture (MM), Max-Sum-Mixture (MSM), Sum-Mixture (SM), the Hessian Sum-Mixture (HSM), and NLS-Hessian Sum-Mixture (NLS-HSM) are all evaluated on two simulated experiments and one real dataset. The HSM approach uses a modified solver that allows specifying the Hessian, while NLS-HSM uses the NLS compatibility method outlined in Sec. VI. The considered simulation examples are similar to the ones provided in [11], and the real-world experiment is a SLAM problem with unknown data associations. In some of the tested examples, particularly the 2D toy example case, the Sum-Mixture approach fails when using standard Gauss-Newton due to a rank-deficient Hessian. To provide a fair comparison, Levenberg-Marquardt [17] (LM) is used instead of Gauss-Newton. LM adds a multiple of the identity to the approximate Hessian, written as $\mbf{H}^{\text{LM}}=\mbf{H}+\lambda\mbf{I}$, where $\lambda$ is a damping constant for the LM method, $\mbf{H}$ is the original Hessian approximation given by (6) for the Max-Mixture, Sum-Mixture, and Max-Sum-Mixture, and by (31) for the Hessian Sum Mixture approach. This ensures that the Hessian approximation is always full rank. The LM damping constant $\lambda$ is chosen as described in [17].
The metrics used to evaluate the quality of the obtained solution are root-mean-squared error (RMSE), defined as
RMSE $=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left\|\hat{\mbf{x}}_{i}\ominus\mbf{x}_{i}\right\|_{2}^{2}},$ (42)
where $\hat{\mbf{x}}_{i}$ and $\mbf{x}_{i}$ are, respectively, the estimated and ground truth states at timestep $i$. The $\ominus$ operator represents a generalized “subtraction” operator, required as some examples considered involve states defined on Lie groups. Following the definitions in [20], a left perturbation is used for all Lie group operations in this paper. Additionally, for states that involve rotations, the difference between the states is split into a rotational and translational component, yielding a rotational RMSE (deg) and a translational RMSE (m).
Consistency, the ability of the estimator to characterize its error uncertainty, is measured using average normalized estimation error squared (ANEES) [21]. A perfectly consistent estimator has ANEES equal to one. The main drawback of this metric is that it characterizes the posterior state belief as Gaussian, which is never completely true in practice.
Lastly, the average number of iterations is used to characterize how quickly the estimator converges to the solution. A smaller number of iterations indicates better descent directions in (2), which is expected if a better Hessian approximation is used. The solver exits once the step size of the LM algorithm is below $1\times 10^{-8}$ or a cap of 200 iterations is reached.
@@ VII-A Toy Example
Method RMSE (m) Iterations Succ. Rate [%]
1D MM 4.12e-01 1.43 25.6
SM 1.26e-02 26.15 99.3
MSM 1.27e-02 18.77 99.3
${}^{\dagger}$ 1.30e-02 7.15 99.2
NLS ${}^{\dagger}$ 1.30e-02 7.16 99.2
2D MM 6.64e-01 1.57 31.9
SM 5.38e-02 28.00 97.6
MSM 4.71e-02 13.03 97.9
${}^{\dagger}$ 4.81e-02 7.10 97.9
NLS ${}^{\dagger}$ 4.81e-02 7.10 97.9
@@
The toy example experiments in this section use four mixture components, such that ${n_{k}}=4$. To quantify the performance of all algorithms considered, Monte-Carlo trials are run, where in each trial, random mixture parameters are generated and run for many different initial conditions. Both 1D and 2D examples are considered, where in the 1D case, the design variable is a scalar, and in the 2D case, $\mbf{x}\in\mathbb{R}^{2}$. In all trials, the first mixture component is chosen with weight $w_{1}\sim\mathcal{U}(0.2,0.8)$, where $\mathcal{U}(a,b)$ denotes the uniform distribution over the interval $[a,b]$. The weights for the other mixture components component are chosen as $w_{k}=(1-w_{1})/({n_{k}}-1)$. The mean for the first mixture component is set to $\check{\mbf{x}}_{1}=\mbf{0}$, while each component of the mean of the other mixture components is chosen as $\check{x}_{2}^{i}\sim\mathcal{U}(-2,2)$. The covariance for the first component is set to $\mbf{R}_{1}=\sigma_{1}^{2}\mbf{1}$, where $\sigma_{1}^{2}\sim\mathcal{U}(0.4,1)$. The second component covariance is set to a multiple of the first component such that $\mbf{R}_{2}=m\mbf{R}_{1}$, where $m\sim\mathcal{U}(4,10)$. For each generated set of mixture parameters, the optimization is run from 100 different initial points, chosen uniformly on a grid on $[-4,4]$ for each axis in the problem dimension. Additionally, 1000 different mixture parameters are generated, resulting in a total of 100, 000 Monte-Carlo trials for both the 1D example and 2D example. The ground truth is determined by sampling on a grid combined with local optimization to find the global optimum. The result is marked as converged if it converges to the ground truth within a given threshold, set to $0.01$ meters for all experiments. The results for both the 1D and 2D Monte-Carlo experiments are shown in Table I. In all tables, the proposed method is denoted by ${}^{\dagger}$. The posterior in this case is extremely non-Gaussian, thus the ANEES is not reported. All methods achieve similar error values, with Sum-Mixture presenting a small improvement over the other methods. Note however that this is due to its interaction with LM, since for the 2D case particularly, the Gauss-Newton version of Sum-Mixture results in a non-invertible Hessian approximation and the method does not work. Nevertheless, the HSM approach still achieves a lower number of iterations than SM or MSM. The Max-Mixture approach achieves the fewest iterations but has a very low success rate.
@@ VII-B Point-Set Registration
Method RMSE (deg) RMSE (m) ANEES Iter.
2D MM 1.57e+00 3.27e-02 3.10e+00 13.77
SM 1.12e+00 1.97e-02 5.10e-02 56.81
MSM 1.11e+00 1.92e-02 1.60e+00 26.33
${}^{\dagger}$ 1.11e+00 1.92e-02 1.43e+00 23.45
NLS ${}^{\dagger}$ 1.11e+00 1.92e-02 1.43e+00 23.45
3D MM 1.12e+00 2.93e-02 1.82e+00 14.74
SM 9.57e-01 2.15e-02 2.76e-02 70.21
MSM 9.59e-01 2.18e-02 1.35e+00 22.93
${}^{\dagger}$ 9.59e-01 2.18e-02 1.27e+00 21.34
NLS ${}^{\dagger}$ 9.59e-01 2.18e-02 1.27e+00 21.34
The point-set registration example follows the structure of the example in [11], which itself builds on the formulation of [14]. Given an uncertain source set of points =$\mathcal{M}=\left\{{\mbf{m}}_{i}\right\}_{i=1}^{M}$, as well as an uncertain reference set of points ${\mathcal{T}}=\left\{{\mbf{p}}_{j}\right\}_{j=1}^{F}$ with ${\mbf{m}}_{i},{\mbf{p}}_{j}\in\mathbb{R}^{{N}}$, the point-set registration task seeks to find the rigid-body transformation $\mbf{T}_{{t}{s}}^{\star}\in{SE({N})}$ that best aligns ${\mathcal{M}}$ to ${\mathcal{T}}$, where $N$ is either $2$ or $3$. The problem setup is depicted in Figure 2. The transformation is found by forming point-to-point residuals, $\mbf{e}_{ij}$,
@@
where $\boldsymbol{\Sigma}_{m}$, $\boldsymbol{\Sigma}_{f}$ are the noise covariances on point measurements of points in ${\mathcal{M}}$ and ${\mathcal{T}}$, respectively. Similarly to [11], to incorporate data association ambiguity, the likelihood of a measurement $\mbf{m}_{i}$ is given by the Gaussian mixture
$p(\mbf{m}_{i}|\mbf{C}_{ts},\mbf{r}_{t}^{st},{\mathcal{T}})$ $=\sum_{j=1}^{M}w_{j}\mathcal{N}\left(\mbf{e}_{ij}(\mbf{p}_{j},\mbf{m}_{i},\mbf{C}_{ts},\mbf{r}_{t}^{st});\mbf{0},\mbf{R}_{ij}\right),$
@@
with uniform weights $w_{j}=\frac{1}{M}$. 100 different landmark configurations are generated, with 100 different noisy point-cloud pairs generated for each. The experiments are run for a planar 2D case where $N=2$, and a 3D case where $N=3$. For the 2D (3D) case, each landmark configuration is simulated by first generating 15 (20) landmarks where each position component is generated according to $\mathcal{U}\sim(-5,5)$, followed by duplicating $30\%$ of them 4 times in a Gaussian with $0.1\mbf{1}$ covariance around the original. 100 measurement pairs are then generated for each landmark. The uncorrupted measurements are generated using the ground truth transformation $\mbf{T}_{{t}{s}}$. The groundtruth transformation itself is randomly sampled as $\mbf{T}_{{t}{s}}=\exp(\boldsymbol{\xi}_{{t}{s}}^{\wedge})$, where $(\cdot)^{\wedge}:\mathbb{R}^{m}\to\mathfrak{g}$, with $m=3$ for $SE(2)$ and $m=6$ for $SE(3)$, and $\exp\left(\cdot\right):\mathfrak{g}\to G$ is the corresponding exponential map for $SE(2)$ or $SE(3)$. The Lie group and its corresponding Lie algebra are denoted $G$ and $\mathfrak{g}$ respectively. In all presented experiments, $\boldsymbol{\xi}_{{t}{s}}^{\mathsf{T}}=\begin{bmatrix}\boldsymbol{\xi}^{\phi^{\mathsf{T}}}&\boldsymbol{\xi}^{\textrm{r}^{\mathsf{T}}}\end{bmatrix}$ with $\boldsymbol{\xi}^{\phi}\sim\mathcal{U}(-\frac{15}{180},\frac{15}{180})$, and $\boldsymbol{\xi}^{\textrm{r}}\sim\mathcal{U}(-0.5,0.5)$. The source and target point covariances $\boldsymbol{\Sigma}_{m},\boldsymbol{\Sigma}_{f}$ are each generated randomly as $\boldsymbol{\Sigma}=\mbf{C}\mbf{D}\mbf{C}^{\mathsf{T}}$, where $\mbf{D}$ is a diagonal matrix with entries generated from $\mathcal{U}(0.1,0.6)$ and $\mbf{C}\in SO(N)$ is a random direction-cosine-matrix generated as $\mbf{C}=\exp(\boldsymbol{\xi}^{\wedge})$. Each entry in $\boldsymbol{\xi}$ is generated as $\xi_{i}\sim\mathcal{U}(-\pi,\pi)$, and the $(\cdot)^{\wedge}$ and $\exp(\cdot)$ operators are overloaded for $SO(N)$. Since the state is now defined on a Lie group, the Jacobian and Hessian of the loss with respect to the state are now replaced with their Lie group counterparts [20],
$\frac{DJ}{D\mathcal{X}}$ $\triangleq\lim_{\boldsymbol{\tau}\rightarrow\mbf{0}}\frac{J(\mathcal{X}\oplus\boldsymbol{\tau})\ominus J(\mathcal{X})}{\boldsymbol{\tau}},$ (46)
$\frac{D^{2}J}{D\mathcal{X}^{2}}$ $\triangleq\lim_{\boldsymbol{\tau}\rightarrow\mbf{0}}\frac{\left.\frac{DJ}{D\mathcal{X}}\right|_{\mathcal{X}\oplus\boldsymbol{\tau}}\ominus\left.\frac{DJ}{D\mathcal{X}}\right|_{\mathcal{X}}}{\boldsymbol{\tau}},$ (47)
where $\oplus$ is the generalized “addition” operator for Lie groups defined analagously to the $\ominus$ operator [20]. The results are presented in Table II. In this case, the proposed HSM approach slightly improves ANEES, as well as improving the number of iterations to convergence. The RMSE is similar for all approaches, with the exception of Max-Mixture, which has a higher RMSE, although it converges the fastest.
Figure 2: 2D Point-Set Registration Setup. Using the uncertain point positions and measurements, the optimization seeks the transformation from the source frame to the reference frame.

### v2

@@ V Proposed Approach: Hessian-Sum-Mixture
The proposed Hessian-Sum-Mixture Hessian approximation is derived by setting the second derivatives of the component errors to zero, and using the chain rule to take into account the LogSumExp nonlinearity and derive the cost Hessian.
@@ VI Nonlinear Least Squares Compatibility
and substituting into the GN Hessian approximation (5) results in exactly the Hessian-Sum-Mixture Hessian (26) as well as the loss Jacobian (25). However, the resulting cost function $\left\|\mbf{e}_{\textrm{solver},1}\right\|_{2}^{2}$ is not of the same form as (10), which can cause issues with methods that use the cost function to guide the optimization such as Levenberg-Marquardt (LM) [madsen2004methods]. To rectify this, an additional error term is required to make the loss into the same form as (10). The proposed error and Jacobian to be input into the solver are thus
$\mbf{e}_{\textrm{solver}}^{\mathsf{T}}$ $=\begin{bmatrix}\mbf{e}_{\textrm{solver},1}^{\mathsf{T}}&\sqrt{{{\color[rgb]{0,0,0}2(}}{\gamma_{\text{{{HSM}}}}}+\Delta J{{\color[rgb]{0,0,0})}}}\end{bmatrix},$ (33)
$\mbf{J}_{\textrm{solver}}^{\mathsf{T}}$ $=\begin{bmatrix}\mbf{J}_{\textrm{solver, 1}}^{\mathsf{T}}&0\end{bmatrix},$ (34)
where $\Delta J$ is the difference between the desired loss $J_{\textrm{GMM}}$ (10) and the squared norm of $\mbf{e}_{\textrm{solver},1}$,
$\Delta J=J_{\textrm{GMM}}-{{\color[rgb]{0,0,0}\frac{1}{2}}}\left\|\mbf{e}_{\textrm{solver},1}\right\|_{2}^{2},$ (35)
and ${\gamma_{\text{{{HSM}}}}}$ is a normalization constant that ensures positivity of the square root in (33). The evaluated cost $\frac{1}{2}\left\|\mbf{e}_{\textrm{solver}}\right\|_{2}^{2}$ is thus ${{\color[rgb]{0,0,0}\frac{1}{2}}}\left\|\mbf{e}_{\textrm{solver}}\right\|_{2}^{2}=J_{\textrm{GMM}}+{\gamma_{\text{{{HSM}}}}}$. Note that since $\mbf{J}_{\textrm{solver}}$ is not the true Jacobian of the error $\mbf{e}_{\textrm{solver}}$, caution must be taken if using automatic differentiation. The normalization constant ${\gamma_{\text{{{HSM}}}}}$ is taken to be the lower bound on $\Delta J$ to ensure positivity of the square root argument. By taking the logarithm and exponent of the second term in (35), and using the properties of exponents, (35) becomes
$\Delta J$ $=-\log\sum_{k=1}^{n_{k}}\left(\alpha_{k}\exp\left(S_{k}\right)\right),$ (36)
$S_{k}$ $=\sum_{j=1}^{n_{k}}\frac{\alpha_{j}\exp\left(-f_{j}\right)}{\sum_{i=1}^{n_{k}}\alpha_{i}\exp\left(-f_{i}\right)}\left(f_{j}-f_{k}\right),$ (37)
with $f_{k}=\frac{1}{2}\mbf{e}_{k}^{\mathsf{T}}\mbf{e}_{k}$. A sequence of algebraic manipulations and using the fact that $t\exp(-t)\leq\exp(-1)$ yields $S_{k}\leq\frac{1}{\alpha_{k}}\sum_{j=1}^{n_{k}}\alpha_{j}$, and thus
$\Delta J\geq-\log\sum_{k=1}^{n_{k}}\left(\alpha_{k}\exp\left(\frac{1}{\alpha_{k}}\sum_{j=1}^{n_{k}}\alpha_{j}\right)\right).$ (38)
The negative of the lower bound (38) is used as ${\gamma_{\text{{{HSM}}}}}$.
@@ VII Simulation and Experiments
To show the benefits of the proposed approach, the performance of the Max-Mixture (MM), Max-Sum-Mixture (MSM), Sum-Mixture (SM), the Hessian Sum-Mixture (HSM), are all evaluated on two simulated experiments and one real dataset. A benefit of HSM is in treating all of the components in the same way, unlike MM or MSM that privilege the dominant component. The simulated examples thus focus on cases where components overlap. The considered simulation examples are similar to the ones provided in [pfeiferAdvancingMixtureModels2021], and the real-world experiment is a SLAM problem with unknown data associations. In some of the tested examples, particularly the 2D toy example case, the Sum-Mixture approach fails when using standard Gauss-Newton due to a rank-deficient Hessian. To provide a fair comparison, Levenberg-Marquardt [madsen2004methods] (LM) is used instead of Gauss-Newton for all approaches. LM adds a multiple of the identity to the approximate Hessian, written as $\mbf{H}^{\text{LM}}=\mbf{H}+\lambda\mbf{I}$, where $\lambda$ is a damping constant for the LM method, $\mbf{H}$ is the original Hessian approximation given by (5) for the Max-Mixture, Sum-Mixture, and Max-Sum-Mixture, and by (30) for the Hessian Sum Mixture approach. This ensures that the Hessian approximation is always full rank. The LM damping constant $\lambda$ is set as described in [madsen2004methods].
The metrics used to evaluate the quality of the obtained solution are root-mean-squared error (RMSE), defined as $\text{RMSE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left\|\hat{\mbf{x}}_{i}\ominus\mbf{x}_{i}\right\|_{2}^{2}}$, where $\hat{\mbf{x}}_{i}$ and $\mbf{x}_{i}$ are, respectively, the estimated and ground truth states at timestep $i$. The $\ominus$ operator represents a generalized “subtraction” operator, required as some examples considered involve states defined on Lie groups. Following the definitions in [sola2021micro], a left perturbation is used for all Lie group operations in this paper. Additionally, for states that involve rotations, the difference between the states is split into a rotational and translational component, yielding a rotational RMSE (deg) and a translational RMSE (m).
Consistency, the ability of the estimator to characterize its error uncertainty, is measured using average normalized estimation error squared (ANEES) [Li2002]. A perfectly consistent estimator has ANEES equal to one. The main drawback of this metric is that it characterizes the posterior state belief as Gaussian, which is never completely true in practice.
Lastly, the average number of iterations and runtime are used to characterize how quickly the estimator converges to the solution. A smaller number of iterations indicates better descent directions in (1), which is expected if a better Hessian approximation is used. The solver exits once the step size of the LM algorithm is below $1\times 10^{-8}$ or a cap of 200 iterations is reached.
@@ VII-A Toy Example
Dims. Method RMSE Avg Time Succ.
(m) Iter. (s) Rate [%]
1D MM 4.07e-01 1.4 4.24e-03 24.1
SM 1.84e-02 26.1 4.12e-02 98.9
MSM 1.66e-02 18.6 3.10e-02 99.0
HSM 1.67e-02 8.8 1.50e-02 99.0
2D MM 7.16e-01 1.5 4.42e-03 30.4
SM 5.30e-02 27.9 4.61e-02 97.5
MSM 4.83e-02 12.9 2.31e-02 97.8
HSM 4.89e-02 9.1 1.57e-02 97.8
@@
The toy example experiments in this section use four mixture components, such that ${n_{k}}=4$. To quantify the performance of all algorithms considered, Monte-Carlo trials are run, where in each trial, random mixture parameters are generated and run for many different initial conditions. Both 1D and 2D examples are considered, where in the 1D case, the design variable is a scalar, and in the 2D case, $\mbf{x}\in\mathbb{R}^{2}$. In all trials, the first mixture component is chosen with weight $w_{1}\sim\mathcal{U}(0.2,0.8)$, where $\mathcal{U}(a,b)$ denotes the uniform distribution over the interval $[a,b]$. The weights for the other mixture components component are chosen as $w_{k}=(1-w_{1})/({n_{k}}-1)$. The mean for the first mixture component is set to $\check{\mbf{x}}_{1}=\mbf{0}$, while each component of the mean of the other mixture components is chosen as $\check{x}_{2}^{i}\sim\mathcal{U}(-2,2)$. The covariance for the first component is set to $\mbf{R}_{1}=\sigma_{1}^{2}\mbf{1}$, where $\sigma_{1}^{2}\sim\mathcal{U}(0.4,1)$. The second component covariance is set to a multiple of the first component such that $\mbf{R}_{2}=m\mbf{R}_{1}$, where $m\sim\mathcal{U}(4,10)$. For each generated set of mixture parameters, the optimization is run from 100 different initial points, chosen uniformly on a grid on $[-4,4]$ for each axis in the problem dimension. Additionally, 1000 different mixture parameters are generated, resulting in a total of 100, 000 Monte-Carlo trials for both the 1D example and 2D example. The ground truth is determined by sampling on a grid combined with local optimization to find the global optimum. The result is marked as successful if it converges to the ground truth within a given threshold, set to $0.01$ meters for all experiments. The results for both the 1D and 2D Monte-Carlo experiments are shown in Table I. In all tables, the proposed method is denoted by . The posterior in this case is extremely non-Gaussian, thus the ANEES is not reported. All methods achieve similar error values, with Sum-Mixture presenting a small improvement over the other methods. Note however that this is due to its interaction with LM, since for the 2D case particularly, the Gauss-Newton version of Sum-Mixture results in a non-invertible Hessian approximation and the method does not work. Nevertheless, the HSM approach still achieves a lower number of iterations than SM or MSM. The Max-Mixture approach achieves the fewest iterations but has a very low success rate.
@@ VII-B Point-Set Registration
Dims. Method RMSE RMSE ANEES Avg Time
(deg) (m) Iter. (s)
2D MM 1.59 0.03 2.92 13.54 1.75
SM 1.15 0.02 0.05 55.05 9.50
MSM 1.15 0.02 1.61 25.79 4.27
HSM 1.15 0.02 1.44 22.56 3.66
3D MM 1.10 0.03 1.81 14.71 3.06
SM 0.95 0.02 0.03 71.08 20.53
MSM 0.95 0.02 1.34 22.81 6.02
HSM 0.95 0.02 1.26 21.24 5.51
The point-set registration example follows the structure of the example in [pfeiferAdvancingMixtureModels2021], which itself builds on the formulation of [Jian2011]. Given an uncertain source set of points =$\mathcal{M}=\left\{{\mbf{m}}_{i}\right\}_{i=1}^{M}$, as well as an uncertain reference set of points ${\mathcal{T}}=\left\{{\mbf{p}}_{j}\right\}_{j=1}^{F}$ with ${\mbf{m}}_{i},{\mbf{p}}_{j}\in\mathbb{R}^{{N}}$, the point-set registration task seeks to find the rigid-body transformation $\mbf{T}_{{t}{s}}^{\star}\in{SE({N})}$ that best aligns ${\mathcal{M}}$ to ${\mathcal{T}}$, where $N$ is either $2$ or $3$. The transformation is found by forming point-to-point residuals, $\mbf{e}_{ij}$,
@@
where $\boldsymbol{\Sigma}_{m}$, $\boldsymbol{\Sigma}_{f}$ are the noise covariances on point measurements of points in ${\mathcal{M}}$ and ${\mathcal{T}}$, respectively. Similarly to [pfeiferAdvancingMixtureModels2021], to incorporate data association ambiguity, the likelihood of a measurement $\mbf{m}_{i}$ is given by the Gaussian mixture
$p(\mbf{m}_{i}|\mbf{C}_{ts},\mbf{r}_{t}^{st},{\mathcal{T}})$ $=\sum_{j=1}^{M}w_{j}\mathcal{N}\left(\mbf{e}_{ij}(\mbf{p}_{j},\mbf{m}_{i},\mbf{C}_{ts},\mbf{r}_{t}^{st});\mbf{0},\mbf{R}_{ij}\right),$ (42)
@@
with uniform weights $w_{j}=\frac{1}{M}$. 100 different landmark configurations are generated, with 100 different noisy point-cloud pairs generated for each. The experiments are run for a planar 2D case where $N=2$, and a 3D case where $N=3$. For the 2D (3D) case, each landmark configuration is simulated by first generating 15 (20) landmarks where each position component is generated according to $\mathcal{U}\sim(-5,5)$, followed by duplicating $30\%$ of them 4 times in a Gaussian with $0.1\mbf{1}$ covariance around the original. 100 measurement pairs are then generated for each landmark. The uncorrupted measurements are generated using the ground truth transformation $\mbf{T}_{{t}{s}}$. The groundtruth transformation itself is randomly sampled as $\mbf{T}_{{t}{s}}=\exp(\boldsymbol{\xi}_{{t}{s}}^{\wedge})$, where $(\cdot)^{\wedge}:\mathbb{R}^{m}\to\mathfrak{g}$, with $m=3$ for $SE(2)$ and $m=6$ for $SE(3)$, and $\exp\left(\cdot\right):\mathfrak{g}\to G$ is the corresponding exponential map for $SE(2)$ or $SE(3)$. The Lie group and its corresponding Lie algebra are denoted $G$ and $\mathfrak{g}$ respectively. In all presented experiments, $\boldsymbol{\xi}_{{t}{s}}^{\mathsf{T}}=\begin{bmatrix}\boldsymbol{\xi}^{\phi^{\mathsf{T}}}&\boldsymbol{\xi}^{\textrm{r}^{\mathsf{T}}}\end{bmatrix}$ with $\boldsymbol{\xi}^{\phi}\sim\mathcal{U}(-\frac{15}{180},\frac{15}{180})$, and $\boldsymbol{\xi}^{\textrm{r}}\sim\mathcal{U}(-0.5,0.5)$. The source and target point covariances $\boldsymbol{\Sigma}_{m},\boldsymbol{\Sigma}_{f}$ are each generated randomly as $\boldsymbol{\Sigma}=\mbf{C}\mbf{D}\mbf{C}^{\mathsf{T}}$, where $\mbf{D}$ is a diagonal matrix with entries generated from $\mathcal{U}(0.1,0.6)$ and $\mbf{C}\in SO(N)$ is a random direction-cosine-matrix generated as $\mbf{C}=\exp(\boldsymbol{\xi}^{\wedge})$. Each entry in $\boldsymbol{\xi}$ is generated as $\xi_{i}\sim\mathcal{U}(-\pi,\pi)$, and the $(\cdot)^{\wedge}$ and $\exp(\cdot)$ operators are overloaded for $SO(N)$. Since the state is now defined on a Lie group, the Jacobian and Hessian of the loss with respect to the state are now replaced with their Lie group counterparts [sola2021micro]. The results are presented in Table II. In this case, the proposed HSM approach slightly improves ANEES. This is expected since the Hessian approximation influences the resulting Laplace approximation and uncertainty characterization. The RMSE is similar for all approaches, with the exception of Max-Mixture, which has a higher RMSE, although it converges the fastest.